3. In class we analyzed data on whether taller US presidential candidate won the election. Analyze the data for 1932-2012 below (note: Mitt Romney is 1 inch taller than Barrack Obama). HeightFrequency taller shorter 15

(a) by hand

construct 95% confidence interval for the proportion that taller US presidential candidates won the election based on exact binomial distribution.

(b) pi is the probability of success

Find and plot the binomial likelihood function over the space of potential values for π.
Answer #1

a)

p^= X/n = 15/20 = 0.75

k=0 n. k=x

Exact Confidence Interval around Proportion:

to

(0.509,0.9134)

b)

let tall onatt ις 0LS

File History ResizeWindows R R Console R Graphics Device 2 (ACTIVE) p←c (0:10) > p-p/10 121 0.0 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9 1.0 1 .000000e+00 5.904900e-1.073742e-11 2.411621e-09 8.349416e-03 O o 111 o.o00000e+00 >plot (p,) > P-p/100 61 9.536743e-07 4.814694e-06 1.153657e-05 1. 125900e-05 2.058911e-06 plot (.p p-c (D:100) plot (P. y) 0.0 0.2 0.4 06 08 1.0